# CURATe: Benchmarking Personalised Alignment of Conversational AI Assistants

### v1

@@ A.2.3 Example completions
A.3 Experimental setup
A.3.1 Model configurations
All models were configured with a temperature of 0.7 and 100 max new tokens, which was ample for this task. This limit helped was to reduce runtime and avoid unnecessary expenditure.
@@ A.3.1 Model configurations
A.4 Additional visualisations
For additional reference, we included pairwise comparisons of all the different ablations (Figures 1091211) , a heatmap distribution of ambiguity factors across different categories and scenarios (Figure  15), as well as visualisations showing the mean percentage of ambiguous results for each model across scenarios (Figure 14, and the percentage that each category contributed to ambiguous results per scenario, respectively (Figure 13).

### v2

@@ B.2.8 Example completions
B.3 Experimental setup
B.3.1 Model configurations
All evaluated models have a temperature of 0.7 and a maximum output of 100 tokens (apart from the evaluator model, which had 130), which was ample for this task. This limit helped was to reduce runtime and avoid unnecessary expenditure. Each conversation begins with the instruction: “You are an AI assistant. Keep your answers brief and don’t ask questions.” [^9] The specification of the model being an assistant was to ensure that it understood its role in the conversation history, as prior interaction turns were phrased as “User:” and “Assistant:” respectively. The specification to keep answers brief was to ensure that important parts of the response were not cut off with the token limit, and avoiding questions helped limit token use further.
@@ B.3.1 Model configurations
B.4 Additional visualisations
For additional reference, we included pairwise comparisons of all the different ablations (Figures 11101312) , a heatmap distribution of ambiguity factors across different categories and scenarios (Figure  16), as well as visualisations showing the mean percentage of ambiguous results for each model across scenarios (Figure 15, and the percentage that each category contributed to ambiguous results per scenario, respectively (Figure 14).